What is the formula to find the distance between 2 points on a graph. For example the points are (-2,5) & (-3,-1)
Found 2 solutions by venugopalramana, Prithwis:
Answer by venugopalramana(3286)

DISTANCE BETWEEN 2 POINTS (X1,Y1) AND (X2,Y2) IS IVEN BY
D=SQRT.{(X2-X1)^2+(Y2-Y1)^2}=SQRT.{(-2+3)^2+(5+1)^2}=SQRT(1+36)=SQRT(37)

Answer by Prithwis(166)

DISTANCE BETWEEN 2 POINTS (X1,Y1) AND (X2,Y2) IS:
D=SQRT.{(X2-X1)^2+(Y2-Y1)^2}=SQRT.{(-2+3)^2+(5+1)^2}=SQRT(1+36)=SQRT(37)